OpenStudy (anonymous):

Really i need help fast! Solve the equation for y. 4x – 5y = 9 A. y = -4x - 9 B. y = -5/4x - 5/9 C. y = 4/5x - 9/5 D. y = 5/4x + 9/4

OpenStudy (igreen):

4x - 5y = 9 Subtract 4x to both sides: -5y = -4x - 9 y = -4/5x - 9/5

OpenStudy (anonymous):

The last one wasnt a choice

OpenStudy (anonymous):

ok

OpenStudy (igreen):

Sorry..I forgot to put 9/5 but I put 9 instead...

OpenStudy (anonymous):

the first one is the answer

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thats not an answer look again @igreen

OpenStudy (igreen):

Yeah, it's Option C..my bad.. -4 / -5 = 4/5.

OpenStudy (anonymous):

whaaaaaaaaaaaaatttt? im confused now where is the question

OpenStudy (igreen):

@ItsMijy-MijyPlays[YouTube][Minecraft]